OpenStudy (anonymous):

What is $\left(1 \frac{1}{2}\right)^4$ expressed as a mixed number?

OpenStudy (anonymous):



OpenStudy (zarkon):

I would write the 1 and 1/2 as 3/2 and then raise to the 4th power. then convert back to a mixed number

OpenStudy (joannablackwelder):

Do you know how to simplify \[\left( \frac{ 3 }{ 2 } \right)^{4}\]?

OpenStudy (zarkon):

\[\Large\left(\frac{3}{2}\right)^4=\frac{3^4}{2^4}\]